OpenStudy (anonymous):

Jannette can repair a refrigerator in 60 minutes. Liza can repair a refrigerator 20 minutes faster. If they work together, how many minutes does it take them to repair the refrigerator?

OpenStudy (ali2x2):

40 minutes

OpenStudy (ali2x2):

subtract 60 and 20

OpenStudy (anonymous):

There isnt a fraction i have to do first?

OpenStudy (anonymous):

Assuming that 2 people who will repair a refrigerator equally fast will do it twice as fast when working together: Jannete repairs 1/60 refrigerator per minute. Lisa repairs 1/40 refrigerator per minute. Togheter they repair 1/60+1/40= 1/60 + 1,5/60 = 2,5/60=5/120=1/24 refrigerator per minute thus it will take them 24 minutes